# An Investigation of Prompt Variations for Zero-shot LLM-based Rankers

### v1

@@ Abstract
An Investigation of Prompt Variations for Zero-shot LLM-based Rankers
Shuoqi Sun1, Shengyao Zhuang1,2, Shuai Wang1, Guido Zuccon1 1The University of Queensland, 2CSIRO
@@ 1 Introduction
Large Language Models (LLMs) are massively parametrised transformer models that have been undergone training on a large extent of text Zhao et al. (2023). Generative LLMs are capable of generating text in response to some textual input, called prompt Brown et al. (2020), or context. This prompting facility has been used to instruct LLMs to perform specific tasks Brown et al. (2020); Wang et al. (2023); White et al. (2023); Zhuang et al. (2023b); Fan et al. (2023).
@@
In this paper, we investigate the use of LLMs to create zero-shot rankers [^1] We specifically focus on re-rankers, where an initial set of documents is retrieved from the index using a first-stage retriever, and a subset is provided to the re-ranker for producing the final search engine results. For ease of reading, we use rankers, in place of re-rankers throughout the paper. Ma et al. (2023); Sun et al. (2023); Tang et al. (2023); Zhuang et al. (2023a); Qin et al. (2023); Zhuang et al. (2023b). With zero-shot, we mean ranking methods that do not require to be specifically trained for ranking tasks (beyond the pretraining executed to create the backbone LLM). These rankers operate by following the instructions provided in the prompt, which include the query for which the ranking should be produced, and the $k$ documents that should be considered for ranking. Using the LLM, rankers then generate an answer to comply with the ranking instruction. Finally, either the generated answer contains the ranking provided by the method, or the logits of the answer are examined to infer a ranking.
@@
Four families of zero-shot LLM rankers have been proposed: pointwise Zhuang et al. (2023a), pairwise Qin et al. (2023), listwise Ma et al. (2023); Sun et al. (2023); Tang et al. (2023), and setwise Zhuang et al. (2023b). They differ because of the ranking algorithm (or mechanism) implemented in the instructions described in the prompt. For example, in pointwise the LLM is instructed to determine the relevance of a document, while in pairwise the LLM is instructed to determine which of two documents is more relevant. Within each family, one or more methods have been proposed. They typically differ in the backbone LLM, the wording of the prompts, and in how the ranking is formed once the answer is generated by the LLM.
@@
PRP, Qin et al. Passage: {text} Query: {query} Does the passage answer the query?
@@
RankGPT, Sun et al. You are RankGPT, an intelligent assistant that can rank passages based on their relevancy to the query. I will provide you with num passages, each indicated by number identifier []. Rank the passages based on their relevance to query: {query}. {PASSAGES} Search Query: {query}. Rank the num passages above based on their relevance to the search query. The passages should be listed in descending order using identifiers. The most relevant passages should be listed first. The output format should be [] > [], e.g., [1] > [2]. Only response the ranking results, do not say any word or explain.
Table 1: Comparison between the prompts used by PRP Qin et al. (2023) and RankGPT Sun et al. (2023): they do not simply differ in the ranking approach used (PRP: pointwise; RankGPT: listwise), but also in the accessory wording present in the prompt, e.g., the role-playing present in RankGPT (first line of the prompt).
Recent work by Zhuang et al. (2023b) has shown that, once these zero-shot LLM rankers are compared using the same backbone LLM and fixing how the ranking is derived (i.e. generation vs. logits), setwise methods are the most effective. Depending on the dataset, pairwise and listwise methods are similarly effective, with pointwise methods providing lower effectiveness overall. While they did recognise that the use of different backbone LLMs in previous work biased the comparison between methods, they did not identify that the actual prompts used by the different rankers differed not just in terms of the words used to describe the ranking algorithm, i.e. the instruction associated with how scoring should be performed, but also in terms of “accessory wording” used. For example, consider the original prompt used by Qin et al. for their PRP pairwise approach Qin et al. (2023) and that used by Sun et al. for their RankGPT listwise approach Sun et al. (2023) – see Table 1. One can notice that RankGPT’s prompt includes also a “role-playing” component (in the system part of the prompt: “You are RankGPT […]”), absent in PRP.
With this respect, then, we ask ourself: What is the effect of such differences in wording used in the prompts? And, more broadly: Are differences in effectiveness due to the actual ranking algorithm, or they are due to the choice of words used in the prompts? Are differences due to LLM characteristics such as backbone and size?
These are important questions because answering them will give us an understanding of what impacts the effectiveness of LLM rankers, and will influence how methods should be compared in future. We explore these directions of enquiry through a wide array of experiments in which we fix the backbone LLMs of the rankers, and we vary prompts in a controlled manner, categorising different prompt components and investigating their effects.
2 Related Work
2.1 Sensibility of LLMs to Prompt
Previous work has shown that LLMs are sensitive to the prompt formulation. For example, Kim et al. (2023) explored the effect of various prompts and prompting technology on deploying LLMs across Natural Language Generation tasks, showing dependency between prompt and effectiveness. Similar findings were reported by Thomas et al. (2023), who explored different prompt wordings in the context of relevance labelling. Kamruzzaman and Kim (2024) explored how different prompting strategies could influence the presence of social biases in LLM outputs: prompts designed to engage higher cognitive processes could reduce biases compared to simpler prompts. These previous work, among others, have demonstrated that the structure and subtle nuances of prompt phrasing can dramatically impact the performance of the LLMs, across a wide range of tasks and contexts.
2.2 Prompt Optimisation and Self-Optimisers
@@
Strategic prompt design is not only beneficial but necessary to harness the full capabilities of LLMs. Recent studies like that by Yang et al. (2023) have taken this further by investigating LLMs as self-optimisers. These models utilize their generative capabilities to iteratively refine prompts, thereby enhancing their performance on downstream tasks. Guo et al. (2023) integrated evolutionary algorithms with LLMs so that prompts can be dynamically adapted without needing gradient information. Sabbatella et al. (2024)’s method involves a continuous relaxation of the search space, allowing for efficient optimization of prompts; this is suitable for scenarios where only black-box access to LLMs is provided. In our work we do not explore the adaptation of self-optimisers to prompts for zero-shot LLM rankers: but our study motivates pursuing this as a direction for future work.
2.3 Zero-Shot LLM Rankers
@@
We next review generative LLM-based rankers; four ranker families are discussed: pointwise Zhuang et al. (2023a), pairwise Qin et al. (2023), listwise Ma et al. (2023); Sun et al. (2023); Tang et al. (2023), and setwise Zhuang et al. (2023b).
@@
In pointwise, two approaches can be followed: generation Liang et al. (2022); Nogueira et al. (2020) and likelihood Zhuang et al. (2021); Zhuang and Zuccon (2021). In generation, LLMs are prompted with a query-document pair and asked to generate a binary answer (“yes"/“no") to the question of whether the document is relevant to the query, and the likelihood of generating “yes" (extracted from the associated token logits) decides the ranking. In likelihood, a query likelihood model is involved in ranking. The LLM is prompted with the document and asked to produce a relevant query. Then, the raking of documents is based on the likelihood of generating the provided query Sachan et al. (2022), which is obtained from the associated token logits. In our experiments with pointwise we implemented the generation approach, which is more commonly used in previous work.
@@
Pairwise ranking instead compares the relevance of two documents to a single query; these are passed onto the model via the prompt. LLMs are prompted to answer which document is more relevant to the query. The ranking of documents is based on relative relevance Qin et al. (2023).
@@
For listwise ranking a list of documents and a query is passed to the LLM via the prompt. LLMs are driven to generate the document labels of relevant documents in a certain order; the ranking then depends on this Ma et al. (2023); Sun et al. (2023); Pradeep et al. (2023).
@@
In setwise, a set of documents is provided with one query. LLMs are prompted to select the most relevant document, in an iterative way. This allows sorting algorithms to give the ranking results based on the preference to documents, and to stop ranking after $k$ documents have been selected (if more than $k$ were given as input) Zhuang et al. (2023b).
@@
In common across all these rankers is their use of generative LLM backbones in a zero-shot manner, i.e. without the need for further training the LLM for the specific ranking approach.
We finally note a recent, different trend in information retrieval where generative LLMs are used to obtain dense representations of documents and queries (separately, i.e. in a bi-encoder manner), which are then in turn used for dense retrieval: while most require fine-tuning Lee et al. (2024); Wang et al. (2024a, b); BehnamGhader et al. (2024), zero-shot methods are also emerging Zhuang et al. (2024). In this paper we focus on zero-shot LLM-based re-rankers, and thus do not consider these dense retrievers; however we note they are likely affected in the same way from the issues we investigate Zhuang et al. (2024).
@@ 3.1 Prompts Collection and Taxonomy of Ranking Prompt Components
We started by collecting the original prompts used in current zero-shot LLM-based rankers Zhuang et al. (2023a); Qin et al. (2023); Ma et al. (2023); Sun et al. (2023); Tang et al. (2023); Zhuang et al. (2023b); these are reported in Appendix A.
@@
pairwise - Given a query, which of the following two passages is more relevant to the query? - - -
listwise - Rank the {num} passages based on their relevance to the search query. Sort the Passages by their relevance to the Query. I will provide you with {num} passages, each indicated by number identifier []. Rank the passages based on their relevance to query. - -
setwise - Which one is the most relevant to the query. - - - -
@@
pairwise - Output Passage A or Passage B. - - - -
listwise - Sorted Passages = [ The passages should be listed in descending order using identifiers. The most relevant passages should be listed first. The output format should be [] >[], e.g., [1] >[2]. - - -
setwise - Output the passage label of the most relevant passage. Generate the passage label. Generate the passage label that is the most relevant to the query, then explain why you think this passage is the most relevant. - -
@@
Role Playing (RP) All You are RankGPT, an intelligent assistant that can rank passages based on their relevancy to the query. - - - -
After the original prompts were collected, we analysed the prompts to identify high level components that are present in at least one original prompt, along with the associated variants. We also augmented the list of components and variants with other wordings we devised to explore specific categories further, e.g., tone words (see below).
The analysis revealed five components:
@@ 3.2 Building Prompt Variations
QF RP+ TI (Q) + P + TW+ TO RP+ TW+ TO + TI (Q) + P
PF RP+ P + TI (Q) + TW+ TO RP+ TW+ TO + P + TI (Q)
Then, for each template, we consider all possible instantiations. The number of variations of wordings for the task instruction (TI) and output type (OT) components differ across families of rankers – thus giving rise to different number of prompt instantiations across each family. In particular, for pointwise we consider 768 unique prompts; for pairwise 48; for listwise 288; and for setwise 144.
Finally note that we did minor modifications to the original prompts from previous works to improve their consistency: for example, some prompts enclosed the query in quotes, while others appended the query after a colon; other differences included the presence of multiple line breaks. We settled on adapting a unique format for these aspects. These differences in prompt formatting resulted in non statistically significant differences in effectiveness (e.g., Qin et al.’s original prompt on COVID and Llama 3-8B backbone obtained an nDCG@10 of 0.8014, while our adjusted prompt lead to an nDCG@10 of 0.7966).
@@ 3.3 Experiments Settings
To experiment with the LLM rankers and the prompt variations, we setup a two stage ranking pipeline in line with previous work. For the first stage, we use the BM25 implementation from Pyserini Lin et al. (2021) to retrieve the top 100 documents for a query. For the second stage, we use the LLM ranker to re-rank these 100 documents.
As LLMs backbones, we selected instruction-tuned checkpoints from the Flan-T5 family Chung et al. (2024), Mistral 7B Jiang et al. (2023), and Llama3 8B AI@Meta (2024). These are popular and highly-performant open LLMs; we excluded close-source LLMs like GPT4 in our analysis because of the high costs associated with running our experiments on commercial APIs.
For Flan-T5 we considered checkpoints of different sizes: Large (783M), XL (2.85B) and XXL (11.3B). This allowed us to explore the role of backbone size in instruction following ability and ranking. We did not considered extending the experiments to larger sizes of the other backbones, e.g., Llama3 70B, because of the high computational costs associated with doing this systematically.
Flan-T5 models have a maximum input length of 512 tokens. Upon analysis of queries and passages lengths used in our evaluation, methods such as setwise and listwise will exceed the length limits as they have multiple passages in the prompt, and amount all methods we investigated, listwise requires longest input length. Hence, to make comparisons between the different experimental conditions fair,we truncated inputs for all methods and LLM backbones to fit the 512 tokens limit. Based on our calculation of the longest non-query-and-passage prompt instructions, we set the query and passage lengths at 20 and 80 words, respectively.
We used three datasets from the ir_datasets MacAvaney et al. (2021) python library: TREC Deep Learning (DL) 2019 and 2020 (43 and 48 queries respectively), and BeIR TREC COVID (50 queries) Voorhees et al. (2021); Thakur et al. (2021). We performed our analysis using nDCG@10, the primary metric across these datasets.
@@ 4.1 Are there better prompts?
For each family of rankers, we compare the effectiveness of the original prompts with all other prompt variations. Results are displayed in Figure 1: for each family of rankers, the star symbols identify the original prompts, while the boxplot shows the distribution of effectiveness across all prompt variations. In all cases we find better prompts that can achieve statistically higher effectiveness than the original prompts, with the exception of listwise and pairwise on the COVID dataset when the Llama 3 backbone is used. We also identify cases where the original prompt was the worst among those considered for the specific ranking family: this is the case for example for the listwise prompt evaluated on DL19 and DL 20 with the FlanT5-Large backbone. We present the actual nDCG@10 scores of original prompts and the best prompts in Appendix C.

### v2

@@ Abstract
Keywords:
LLM Rankers; Prompt Variations and Optimization
@@ 1 Introduction
Large Language Models (LLMs) are massively parametrised transformer models that have undergone training on a large extent of text [30]. Generative LLMs are capable of generating text in response to textual input (the prompt or context) [3]. This prompting facility has been used to instruct LLMs to perform specific tasks [3, 27, 28, 34, 5, 33].
@@
In this paper, we investigate the use of LLMs to create zero-shot rankers [13, 22, 23, 31, 18, 34]. With zero-shot, we mean ranking methods that are not specifically trained for ranking tasks (i.e. not contrastive fine-tuning). These rankers operate by following the instructions provided in the prompt, which include the query and the $k$ documents that should be considered for ranking. Using the LLM, rankers then generate an answer to comply with the ranking instruction. Finally, either the generated answer contains the ranking provided by the method, or the logits of the answer are examined to infer a ranking.
@@
Four families of zero-shot LLM rankers have been proposed: pointwise [31], pairwise [18], listwise [13, 22, 23], and setwise [34]. They differ because of the ranking algorithm (or mechanism) implemented in the instructions described in the prompt. For example, in pointwise the LLM is instructed to determine the relevance of a document, while in pairwise the LLM is instructed to determine which of two documents is more relevant. Within each family, one or more methods have been proposed. They typically differ in the backbone LLM, the wording of the prompts, and in how the ranking is formed once the answer is generated by the LLM.
Recent work has shown that, once these zero-shot LLM rankers are compared using the same backbone LLM and fixing how the ranking is derived (i.e. generation vs. logits), setwise methods are the most effective [34]. Depending on the dataset, pairwise and listwise methods are similarly effective, with pointwise methods providing lower effectiveness overall. While they did recognise that the use of different backbone LLMs in previous work biased the comparison between methods, they did not identify that the actual prompts used by the different rankers differed not just in terms of the words used to describe the ranking algorithm, i.e. the instruction associated with how scoring should be performed, but also in terms of “accessory wording” used. For example, compare the prompt of the PRP pointwise approach [18] and the one of the RankGPT listwise approach [22] – see Table 1: RankGPT’s prompt includes also a “role-playing” component (in the system part of the prompt: “You are RankGPT […]”), absent in PRP.
With this respect, then, we ask: What is the effect of such differences in wording used in the prompts? And, more broadly: Are differences in effectiveness due to the actual ranking algorithm, or they are due to the choice of words used in the prompts? Are differences due to LLM characteristics such as backbone and size? These are important questions because answering them will give us an understanding of what impacts the effectiveness of LLM rankers, and will influence how methods should be compared in future. We explore these directions of enquiry through a wide array of experiments, for a total of 1,248 prompts, in which we fix the backbone LLMs of the rankers, and we vary prompts in a controlled manner, categorising different prompt components and investigating their effects.
2 Related Work
2.0.1 Sensibility of LLMs to Prompt.
Previous work has shown that LLMs are sensitive to the prompt formulation [9, 25, 10, 8, 2, 15]. For example, Kim et al. [9] explored the effect of various prompts and prompting technology on deploying LLMs across language generation tasks, showing dependency between prompt and effectiveness. Similar findings were reported by Thomas et al. [25] when considering prompt variations in the context of relevance labelling. Kamruzzaman&Kim [8] explored the relation between prompting strategies and social biases in LLM outputs: reduced biases were observed in more complex prompts. The most comprehensive analysis of the effect of prompt variations has been recently provided by Mizrahi et al. across a range of NLP tasks [15]. These previous works, among others, have demonstrated that the structure and subtle nuances of prompt phrasing can dramatically impact the performance of the LLMs, across a wide range of tasks and contexts; however such an analysis does not exist for LLM rankers – our work is novel as it contributes a better understanding of this phenomenon in the specific context of LLM rankers.
Table 1: The PRP prompt (pointwise) [18] differs from that of RankGPT (listwise) [22] also because of accessory wordings (e.g., role-playing in RankGPT).
@@ 2 Related Work
PRP [18] Passage: {text} Query: {query} Does the passage answer the query?
@@
RankGPT [22] You are RankGPT, an intelligent assistant that can rank passages based on their relevancy to the query. I will provide you with num passages, each indicated by number identifier []. Rank the passages based on their relevance to query: {query}. {PASSAGES} Search Query: {query}. Rank the num passages above based on their relevance to the search query. The passages should be listed in descending order using identifiers. The most relevant passages should be listed first. The output format should be [] > [], e.g., [1] > [2]. Only response the ranking results, do not say any word or explain.
2.0.2 Prompt Optimisation and Self-Optimisers.
@@
Strategic prompt design is not only beneficial but necessary to harness the full capabilities of LLMs. Recent studies have taken this further by investigating LLMs as self-optimisers [29, 6, 19]. These models utilize their generative capabilities to iteratively refine prompts, thereby enhancing their performance on downstream tasks. In our work we do not explore the adaptation of self-optimisers to prompts for LLM rankers: but our study motivates pursuing this as a future direction.
2.0.3 Zero-Shot LLM Rankers.
@@
We next review the four families of zero-shot generative LLM-based rankers: pointwise [31], pairwise [18], listwise [13, 22, 23], and setwise [34].
@@
In pointwise, two approaches can be followed: generation [11, 16] and likelihood [32, 35]. In generation, LLMs are prompted with a query-document pair and asked to generate a binary answer (“yes”/“no”) to the question of whether the document is relevant to the query: the likelihood of generating “yes" (extracted from the associated token logits) decides the ranking. In likelihood, a query likelihood model is involved in ranking. The LLM is prompted with the document and asked to produce a relevant query. Then, the ranking of documents is based on the likelihood of generating the provided query [20], which is obtained from the associated token logits. In our experiments with pointwise we implemented the generation approach, which is more commonly used in previous work.
@@
Pairwise methods compare the relevance of two documents to a query, by prompting the LLM to answer which document is more relevant to the query. The ranking of documents is based on the relative preferences [18].
@@
For listwise ranking a list of documents and a query is passed to the LLM via the prompt. LLMs are driven to generate the document labels of relevant documents in a certain order; the ranking then depends on this [13, 22, 17].
@@
In setwise, a set of documents and the query are provided to the LLM, which is prompted to select the most relevant document, in an iterative way. Then sorting algorithms are used to obtain the ranking based on the preferences to documents’ stopping occurs after $k$ documents have been selected (if more than $k$ were given as input) [34].
@@
In common across all these methods is their use of LLM backbones in a zero-shot fashion, i.e. without the need for further training the LLM.
@@ 3.1 Prompts Collection and Taxonomy of Ranking Prompt Components
We started by collecting the original prompts used in current zero-shot LLM-based rankers [31, 18, 13, 22, 23, 34]. After the original prompts were collected, we analysed the prompts to identify high level components that are present in at least one original prompt, along with the associated variants. We also augmented the list of components and variants with other wordings we devised to explore specific categories further, e.g., tone words (see below). The analysis revealed five components:
@@
pairwise - Given a query, which of the following two passages is more relevant to the query? -
listwise - Rank the {num} passages based on their relevance to the search query. Sort the Passages by their relevance to the Query. I will provide you with {num} passages, each indicated by number identifier []. Rank the passages based on their relevance to query. -
setwise - Which one is the most relevant to the query. -
@@
pairwise - Output Passage A or Passage B. -
listwise - Sorted Passages = [ The passages should be listed in descending order using identifiers. The most relevant passages should be listed first. The output format should be [] >[], e.g., [1] >[2]. -
setwise - Output the passage label of the most relevant passage. Generate the passage label. Generate the passage label that is the most relevant to the query, then explain why you think this passage is the most relevant. -
@@
Role Playing (RP) All You are RankGPT, an intelligent assistant that can rank passages based on their relevancy to the query. -
@@ 3.2 Building Prompt Variations
QF RP+ TI (Q) + P + TW+ OT RP+ TW+ OT + TI (Q) + P
PF RP+ P + TI (Q) + TW+ OT RP+ TW+ OT + P + TI (Q)
Then, for each template, we consider all possible instantiations. The number of variations of wordings for the task instruction (TI) and output type (OT) components differ across families of rankers – thus giving rise to different number of prompt instantiations across each family. Overall, our experiments consider 1,248 prompt variations: 768 unique prompts for pointwise; 48 for pairwise; 288 for listwise; and 144 for setwise.
Finally note that we did minor modifications to the original prompts from previous works to improve their consistency: for example, some prompts enclosed the query in quotes, while others appended the query after a colon; other differences included the presence of multiple line breaks. We settled on adapting a unique format for these aspects. From preliminary experiments, we observed that these differences in prompt formatting resulted in non statistically significant differences in effectiveness: e.g., [18]’s original prompt on COVID and Llama 3-8B backbone obtained an nDCG@10 of 0.8014, while our adjusted prompt lead to an nDCG@10 of 0.7966.
@@ 3.3 Experiments Settings
To experiment with the LLM rankers and the prompt variations, we setup a two stage ranking pipeline in line with previous work. For the first stage, we use the BM25 implementation from Pyserini [12] to retrieve the top 100 documents for a query (k1=0.9, b=0.4). For the second stage, we use the LLM ranker to re-rank these 100 documents.
As LLMs backbones, we selected instruction-tuned checkpoints from the Flan-T5 family [4], Mistral 7B [7], and Llama3 8B [1]. These are popular and highly-performant open LLMs; we excluded close-source LLMs like GPT4 in our analysis because of the high costs associated with running our experiments on commercial APIs.
For Flan-T5 we considered checkpoints of different sizes: Large (783M), XL (2.85B) and XXL (11.3B). This allowed us to explore the role of backbone size in instruction following ability and ranking. We did not consider extending the experiments to larger sizes of the other backbones, e.g., Llama3 70B, because of the high computational costs associated with doing this systematically.
We used three datasets from the ir_datasets [14] python library: TREC Deep Learning (DL) 2019 and 2020 (43 and 48 queries respectively), and BeIR TREC COVID (50 queries) [26, 24]. We performed our analysis using nDCG@10, the primary metric across these datasets, and tested statistical significance with a paired, two-tails t-test.
Flan-T5 models are constrained by a maximum input length of 512 tokens. By analysing the query and passage lengths within the evaluation datasets, we found that methodologies like setwise and listwise, which incorporate multiple passages within a single prompt, would exceed this limit, with listwise requiring the longest input length. To ensure equitable comparisons across different experimental setups, we uniformly truncated the inputs for all methods and LLM backbones to adhere to the 512-token constraint. Our determination of truncation lengths was based on a comprehensive analysis of the datasets employed. We calculated the length of the longest non-query-and-passage prompt instruction (pure instruction length) alongside the distribution of query and document lengths. Consequently, we standardized the query length to 20 words and the document length to 80 words. This decision was driven by the need to preserve the integrity of instructional content within the prompts, ensuring that the loss of such information would not compromise experimental consistency across different LLM backbones. In our datasets, while queries do not exceed 20 words, documents can reach up to approximately 120 words, with the majority being about 80 words. Although this truncation might result in some loss of document information, our focus was not on evaluating the absolute ranking performance but on ensuring equitable conditions for each ranker across different backbones. This strategy ensures that, even with the most extensive prompt variations and the shortest model input capabilities, no instructional wording is omitted.
@@ 4.1 Are there better prompts?
For each family of rankers, we compare the effectiveness of the original prompts with all other prompt variations. Results are displayed in Figure 1: for each family of rankers, the star symbols identify the original prompts, while the boxplot shows the distribution of effectiveness across all prompt variations. We found prompts that can achieve higher effectiveness than the original prompts across all cases (with most differences being statistically significant), with the exception of listwise and pairwise on COVID when the Llama 3 backbone is used. For the pairwise approach, only prompts for the FlanT5-XXL model on the COVID dataset show significant differences, likely due to the lower effectiveness variation inherent in the pairwise method. We also identify cases where the original prompt was the worst among those considered for the specific ranking family: for example for the listwise prompt evaluated on DL19 and DL 20 with the FlanT5-Large backbone. We present the actual nDCG@10 scores of original prompts and the best prompts in Table 4.
@@
Table 4: nDCG@10 across rankers, LLMs and datasets. Po: pointwise, Pa: pairwise, Li: listwise, Se: setwise. Original: the original (adapted) prompt. Best: the best prompt in our experiments. For each dataset/backbone – italics: the best performing method that used the original prompt; bold: the best performing method overall. Statistical significance differences (: $p<0.05$, : $p<0.01$) between the original prompt with the best prompt measured with two-tails paired t-test.
Mistral-7B Llama 3-8B
DL19 DL20 COVID DL19 DL20 COVID
Po [18] 0.5851 0.6699 0.5237 0.6486 0.6897 0.7988 0.4176 0.5328 0.3383 0.4950 0.6678 0.7916
Po [13] 0.6254 0.5948 0.7379 0.4808 0.3794 0.7527
Po [31].1 0.6414 0.5951 0.7758 0.3764 0.3088 0.7485
Po [31].2 0.5577 0.4896 0.6967 0.3890 0.3466 0.7242
Pa [18] 0.6263 0.6450 0.6036 0.6059 0.7724 0.7738 0.6738 0.6738 0.6163 0.6340 0.7966 0.8012
Li [13] 0.5863 0.6686 0.5492 0.6348 0.7101 0.7758 0.6364 0.7004 0.6088 0.6619 0.7393 0.8123
Li [22] 0.5402 0.5184 0.6204 0.5926 0.5377 0.6589
Se [34] 0.6408 0.6811 0.5988 0.6256 0.7979 0.8053 0.6654 0.7000 0.6147 0.6486 0.7933 0.8035
FlanT5-Large FlanT5-XL FlanT5-XXL
DL19 DL20 COVID DL19 DL20 COVID DL19 DL20 COVID
Po [18] 0.6180 0.6918 0.5984 0.6266 0.6490 0.7570 0.6157 0.7010 0.6439 0.6727 0.6786 0.7742 0.6289 0.6860 0.6514 0.6733 0.6642 0.7784
Po [13] 0.6483 0.6128 0.6971 0.6690 0.6426 0.7301 0.6563 0.6533 0.7321
Po [31].1 0.6468 0.5698 0.6672 0.6300 0.6444 0.7023 0.6451 0.6274 0.6983
Po [31].2 0.5164 0.4387 0.6206 0.6054 0.5971 0.7095 0.4068 0.3301 0.4335
Pa [18] 0.6677 0.6724 0.6237 0.6382 0.7558 0.7788 0.6845 0.6986 0.6766 0.6823 0.7536 0.7750 0.6915 0.7135 0.6992 0.7126 0.7452 0.7917
Li [13] 0.5465 0.6443 0.5081 0.6125 0.6067 0.7521 0.5688 0.6441 0.5576 0.6300 0.6312 0.7438 0.5915 0.7011 0.5852 0.6897 0.6955 0.7834
Li [22] 0.6199 0.5552 0.7445 0.6129 0.5966 0.6938 0.6920 0.6672 0.7736
Se [34] 0.6535 0.6693 0.6003 0.6525 0.7618 0.7932 0.6859 0.6959 0.6724 0.6855 0.7543 0.7745 0.6864 0.7047 0.6679 0.7036 0.7344 0.7890